OpenStudy (anonymous):

An object moves along a coordinate line, its position at each time t> 0 given by x (t). Find the acceleration at time t0. t(0) = 3

OpenStudy (stormfire1):

Is there any more information to this question?

OpenStudy (stormfire1):

It definitely seems to be missing something :)

OpenStudy (anonymous):

let me3 checl

OpenStudy (stormfire1):

Maybe a picture?

OpenStudy (anonymous):

An object moves along a coordinate line, its position at each time t> 0 given by x (t). Find the acceleration at time t0. https://assessment.casa.uh.edu/Assessment/Images//1019445.gif

OpenStudy (stormfire1):

AHH...much better :) One sec while I look it over

OpenStudy (stormfire1):

Yes...one sec...phone

OpenStudy (stormfire1):

Is this a calc-based course? Just curious

OpenStudy (anonymous):

pre-calc i mean

OpenStudy (stormfire1):

I'm still with you....just checking over my work here

OpenStudy (stormfire1):

You're given: \[x(t)=(t^2-3t)(t^2+3t)\]and \[x(0)=3\]So I think I'd start off by factoring the equation:\[3=(t^2-3t)(t^2+3t)=t^4-9t^2\]\[t^4-9t^2-3=0\]Since velocity is the derivative of position and acceleration is the derivative of velocity, you need to get the second derivative of this equation: \[V(t)=X(t)'=4t^3-18t\]\[A(t)=V(t)'=12t^2-18\] Now that you have the formula for acceleration, you can plug in t = 0 to get -18m/s^2. Hopefully that's one of your choices :)

OpenStudy (anonymous):

lemme check

OpenStudy (anonymous):

the answer choices are a) -54, b) 54, c) 90, d)-14, e)756

OpenStudy (anonymous):

did i mention that t(0) = 3

OpenStudy (stormfire1):

Ok...let me see what I did wrong

OpenStudy (stormfire1):

Ah...yes. I forgot to plug in 3 as you said. If you plug in 3 to the last equation you get 90

OpenStudy (anonymous):

thank you so much.... :)

OpenStudy (stormfire1):

np :)